Wednesday, May 31, 2017
5/30/2017 17:34 PM - Node dalstorage4 lost connectivity. Upon checking the serial console we found a kernel panic occured. 5/30/2017 17:59 PM - All VPS are back onine.
Wednesday, May 24, 2017
We are aware of a performance degradation on dalhighram17. Our engineers are investigating the issue.
Wednesday, May 3, 2017
05/03/2017 06:23 AM CST - We are going to reboot lapure5 to fix the high load average issue on it which was caused by a bug in KernelCare. We utilize KernelCare to do rebootless kernel upgrades but unfortunately a bug in the software is causing lock ups during the patching process. We will update this announcement once all VPS's are back online again. 05/03/2017 07:07 AM CST - All the VPS's...
Wednesday, April 26, 2017
04/26/2017 10:24 CST - We're aware of the network connectivity issue in Dallas and are working on the situation now. We'll update this announcement as we have more information. 04/26/2017 10:30 CST - We have pointed and resolved the issue. It will take around 5 mins before your VPS to come online if it's not already. 04/26/2017 10:54 CST - Everything should be back online now. A full Reason...
Saturday, March 18, 2017
As scheduled, we are currently upgrading Linux Kernel's of the Premium VPS line host nodes. Update: This has been done. Everything is back online.
Monday, March 6, 2017
06 March 2017 14:40 PM - Due to an unexpected hardware issue with ukpure7 VPS node, we are going to perform emergency maintenance on it at 22:00 GMT today, around 1.25 hours from now. Your VPS will be shut down at around 21:50 GMT to prepare the node for the maintenance. Your data and IP address will not change. We expect the server to be back online in around 30 minutes after the...
Sunday, March 5, 2017
We had to upgrade the kernel of all Premium VPS nodes due to the CPU lockups we were getting with the old kernel. We are sorry for the inconvenience.
Thursday, March 2, 2017
02 March 2017 18:49 PM - We are aware that dalstorage4 is down at the moment. We are currently troubleshooting. 02 March 2017 20:12 PM - We were able to boot the server and we are running fsck on the storage at the moment. 02 March 2017 21:36 PM - fsck process has been completed succesfully. We are starting the VPS's on the node now. 02 March 2017 22:01 PM - All the VPS's...
Tuesday, February 28, 2017
Due to issues with Amazon S3, our Amazon SES system which we use to send emails is not working at the moment. Please bare with us until this is resolved. More information: https://status.aws.amazon.com/ Update 2:06 PM: We have switched to another SMTP provider for the time being. UPdate 6:50 PM: We have switched back to Amazon SES.
Thursday, January 26, 2017
26 January 2017, 11:55 - We are investigating the issue within our Dallas network. Please do not open a ticket as this would slow us down at the moment. 26 January 2017, 12:02 - The issue is related to our upstreams from what we understand so far. 26 January 2017, 12:19 - Connectivity has returned. We're awaiting confirmation of stability from the datacenter.